OpenStudy (anonymous):

Use linear regression to find an linear function that best fits this data. (1, 1102) (2,1824) (3,2904) (4,5051) (5,7732) (6,12292)

OpenStudy (lakota1999):

(1,1102

OpenStudy (anonymous):

I need an equation

OpenStudy (lakota1999):

i will only do an equation if u tell me where u live

OpenStudy (anonymous):

wth

OpenStudy (lakota1999):

state

OpenStudy (anonymous):

why?

OpenStudy (anonymous):

@harewood @highgrades02 @Awesome781 @dianaeshun @yeujhkla @pollito___o Can any of you help me? PLEASE

OpenStudy (lakota1999):

(2,1824) (3,2904) (4,5051) (5,7732) (6,12292)1, 1102)

OpenStudy (mathmate):

The regression line is of the form y=ax+b, and the parameters a, b can be found using formulas given in step 2 of the following link: http://www.statisticshowto.com/how-to-find-a-linear-regression-equation/ this assumes that you have the data of the whole population. If your data represent a sample, the resulting regression line is just an approximate one.

OpenStudy (mathmate):

@IloveCharlie By the way, my neighbour's dog's name is Charlie. Your screen name suits me just fine.

OpenStudy (anonymous):

Hey thanks but I still don't get it Can u maybe write it out for me and I will calculate it? PLEASE

OpenStudy (anonymous):

@lakota1999 are you a stalker you creep?

OpenStudy (anonymous):

Lol I think he/she is... @chucho78

OpenStudy (mathmate):

The equation says: Let the regression line be y=ax+b, and n=number f data points then \(\Huge a=\frac{n\Sigma xy-\Sigma x \Sigma y}{n\Sigma x^2 - (\Sigma x)^2}\) \(\huge b=\frac{\Sigma y}{n}-a\frac{\Sigma x}{n}\) (note: I do not use the same notations as in the link.) To calculate \(\Sigma x, \Sigma y, \Sigma x^2, \Sigma xy\) you would be well advised to create a table and add at the bottom of each column, or better still, use Excel.